Beagle Beauties engages in the development, manufacture, and sale of a line of cosmetics designed to make your dog look glamorous. Below you will find selected information necessary to compute some valuation estimates for the firm. Assume the values provided are from year-end 2015. Also assume that the firm’s equity beta is 1.40, the risk-free rate is 2.75 percent, and the market risk premium is 7 percent.

a. What are the sustainable growth rate and required return for Beagle Beauties?

b. Using these values, estimate the current share price of Beagle Beauties stock according to the constant dividend growth model.

c. Using the P/E, P/CF, and P/S ratios, estimate the 2016 share price for Beagle Beauties.Dividends per share Return on equity Book value per share $ 2.04 9. 50% $17.05 Earnings Cash Flow Sales $5.00 2015 value per share Average price multiple Forecasted growth rate 6.60 $25.65 13.10 13 . 48% 9.42 11.41% 2.36 7 . 34%
Answer #1

a) Sustainable growth rate = ROE*retention rate
ROE=9.50%
Retention = 1-(Dividend/EPS)
Retention = 1-(2.04/5.00)=0.592

SGR=9.50*0.592=5.624%

Required return = Rf+(Beta*risk premium)
=2.75+(1.40*7)
=12.55%

b) Price of stock = D1/(r-g)
D1=2.04+(1+SGR)
=2.04*1.05624
r-g=0.1255-0.05624=0.06926
Price=2.04*1.05624/0.06926=$31.11

c)On P/E basis=13.10*5.00*1.1348=$74.33
On P/CF basis= 6.60*9.42*1.1141=$69.27
On P/S basis= 25.65*2.36*1.0734=$64.98